In the weekend, I ran a free swing test on all optics. During this test, LO2 magnets got stuck to the coil because LO2 PIT alignment was very high, making its lower OSEMs almost fully dark and upper OSEMs almost fully bright. Today we realized that LO2 is actually stuck and is not coming off even when we dither PIT alignment. We tried several ways but could not get this off.
Do we have any other method to get magnets off in vaccum?
It will be pretty bad if we try anything related to BHD with LO beam reflecting off a stuck mirror. Does anyone have any suggestions other than venting and fixing the issue?